# Sketch-Plan-Generalize: Continual Few-Shot Learning of Inductively Generalizable Spatial Concepts

### v2

@@ Abstract
Our goal is to enable embodied agents to learn inductively generalize-able spatial concepts, e.g., learning staircase as an inductive composition of towers of increasing height. Given a human demonstration, we seek a learning architecture that infers a succinct program representation that explains the observed instance. Additionally, the approach should generalize inductively to novel structures of different size or complex structures expressed as a hierarchical composition of previously learned concepts. Existing approaches that use code generation capabilities of pre-trained large (visual) language models as well as purely neural models show poor generalization to a-priori unseen complex concepts. Our key insight is to factor inductive concept learning as: (i) Sketch: detecting and inferring a coarse signature of a new concept (ii) Plan: performing MCTS search over grounded action sequences (iii) Generalize: abstracting out grounded plans as inductive programs. Our pipeline facilitates generalization and modular re-use enabling continual concept learning. Our approach combines the benefits of code generation ability of large language models (LLM) along with grounded neural representations, resulting in neuro-symbolic programs that show stronger inductive generalization on the task of constructing complex structures in relation to LLM-only and neural only approaches. Furthermore, we demonstrate reasoning and planning capabilities with learned concepts for embodied instruction following.
@@ 1 Introduction
The ability to learn novel concepts which generalize inductively is one of the hallmarks of human intelligence [mind]. Humans are highly data efficient – observing at a few instances of towers of a certain heights, we can generalize to constructing towers of any height. Further, we interpret increasingly complex concepts as hierarchical composition over simpler ones, e.g., a tower as a sequence of blocks placed on top of each other, or a staircase composed of towers of increasing height.
This paper considers the problem of learning a program representation, from few demonstration, that models the inductive realization of grounded spatial concepts. Learning of such concepts is challenging task due to an expansive space of programs and the need to reason about their physical plausibility. Further, the representation must support inductive generalization over learned concepts as well as express complex hierarchical concepts via modular re-use of concepts learnt previously.
@@
Figure 1: Learning inductive spatial concepts. (Left) A human demonstrates the construction of a row and tower of size three. (Right) The agent learns program representation that enables inductive generalization to novel structures (varied sizes and visual attributes) and expresses complex concepts as hierarchical composition of previously acquired ones. E.g., learning a tower as a sequence of blocks placed one on top of another and a pyramid as rows of decreasing size.
Prior efforts such as [code-as-policies] uses a LLM (Large Language Model) to generate control program for a given task specification but fail to generalize to complex spatial concepts which are difficult to tokenize. Extension of this to VLMs (Vision Language Models), [achiam2023gpt], also fail to generalize when presented with linguistically novel concepts, due to over reliance on prior knowledge and their inability to effectively learn novel concepts from given demonstrations. On the other hand, neural approaches such as  [struct-diffusion], learn from given demonstrations but generalize poorly due to their inability to (a) explicitly model symbolic concept of induction and (b) modularize as well as re-use previously acquired concepts. Approaches such as [li2019practical] train RL-based policies to attain spatial assembly by encoding an inductive spatial prior using GNNs within the policy architecture. However, generalization is still limited and assumes an elicitation of the goal as per object positions, resulting in lack of ability to take in goal description such as “construct a tower of size three”. In summary, we attribute poor generalization of such approaches to an implicit entangling of the following objectives: (i) postulating a high-level program for new concepts, (ii) evaluating plausibility of grounded plans to align with human demonstration of concepts to be learned and (iii) abstracting out a grounded program to facilitate inductive generalization and modular re-use in a continual manner.
In response, this paper introduces an approach that factorizes the concept learning task as: (a) Sketch: Given a language-annotated demonstration of a novel concept, an LLM is used to postulate a function signature. (b) Plan: Refinement of program sketches via MCTS search, rapidly evaluating actions sequences guided by a reward associated with constructing a concept. The search is accelerated by training a neural action predictor that uses the given demonstrations. (c) Generalize: Leveraging the code generation capability of an LLM to distill grounded plans into a program that is inductively generalize-able. This results in a continually evolving library of concepts which can be used to hierarchically learn complex concepts in future. The modular architecture enables continual learning by providing the ability to decide whether the new concept encountered either as a symbolic composition of existing concepts, or, a neural embedding trained via gradient update.
Our experimental evaluation demonstrates accurate learning of simple and complex concepts from few demonstrations for a range of spatial structures. Further, our approach also shows inductive generalization in out-of-distribution settings, significantly over the baselines. We also present deeper insights related to efficiency gains obtained by combining symbolic MCTS with neural action predictor. Furthermore, we show how learned concepts can be grounded in the visual input, enabling a robot to follow natural language instructions referring to a-priori unseen spatial configurations.
@@ 2 Related Works
The ability to ground symbolic knowledge is central to building intelligent embodied agents. The symbol grounding problem is often cast as learning an association between linguistic description of a concept and aspects of the state-action space such as objects [paul2018efficient, ramp], spatial relations  [tellex2011approaching, kim2024lingo], reward functions  [boularias2015grounding], or motion constraints  [howard2014natural]. These works assume the presence of grounded representation for symbolic concepts and only learn associations between language and concepts. In contrast our work jointly learns higher-order concepts composed of simpler concepts along with their grounding in the robot’s state and action space.
Another line of work  [singh2022progprompt, wang2023demo2code, ahn2022can, code-as-policies] leverages prior-knowledge embodied in LLMs to determine grounded robot control programs for a high-level tasks, given a tokenized description of the scene and robot capabilities. Our experiments (reported subsequently) demonstrate that complex spatial relations are challenging for textual descriptions to encode and are difficult to reason over. Hence, we adopt a factored approach that uses a search procedure guided by visual reward, acting as a medium to connect the task sketch generation and program abstraction stages (performed by an LLM).
Works such as [wang2023voyager, jarvis-1] shows lifelong learning of skills by learning to plan high-level tasks through composition of simple skills for simulated agents. Efforts such as  [lotus],  [lifelong-halp] detect when control primitives for a robot are insufficient to perform a task and initiate skill learning, building a library over time. However, such efforts do not model deep inductive use of learned concepts.
Finally, this work draws inspiration from concept learning approaches in the broader AI community. Related works infer latent generative programs to express hand-written digits [lake2015human], object arrangements [ellis2018learning], motion plans [nscl] or goal-directed policies [lpp]. These approaches do not explicitly model inductive concept of modular re-use of previously acquired concepts. Other efforts such as [li2019practical] learn RL-based policies to construct object stacks. Here, the neural policy models locality of object influence via GNNs. However, the model suffers from poor generalization to unseen examples, (eg. tower of larger size) and do not possess a mechanism to re-use previously acquired concepts.
3 Background and Problem Formulation
@@
We consider an embodied agent that uses a visual and depth sensor to observe its environment and can grasp and release objects at specified poses. We represent the robot’s domain as a goal-conditioned MDP $<\mathcal{S},\mathcal{A},\mathcal{T},g,\mathcal{R},\gamma>$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ is the transition function, $g$ is the goal, $R$ is the reward model and $\gamma$ is the discount factor. The agent’s objective is to learn a policy that generates a sequence of actions from an initial state $s_{0}$ to achieve the goal $g$ in response to an instruction $\Lambda$ specifying the intended goal from a human. We assume that the agent possesses a model of semantic relations (e.g., left(), right() etc.) as well as semantic actions such as moving an object by grasping and releasing at a target location. Such modular and composable notions can be acquired from demonstrations via approaches outlined in [nsrm-Kalithasan2022LearningNP, nscl, mao2022pdsketch]. Such notions populate a library of concepts $\mathcal{L}$ available as grounded executable function calls. Following recent efforts [code-as-policies, Huang2022InnerME, ahn2022can, singh2022progprompt] in representing robot control directly as executable programs, we represent action sequence corresponding to a plan as a program consisting of function calls to executable actions and grounded spatial reasoning.
@@
Our goal is to enable a robot to interpret instructions such as “construct a tower with red blocks of height five". Specifically, we aim to learn spatial constructs like a tower that requires sequential actions that repeatedly places a block on top of a previously constructed assembly; a process akin to induction. Given a few demonstrations of a constructing a spatial assembly, $\mathcal{D}$, each consisting of natural language description $\Lambda$ (“construct a tower of red blocks of size five") and a sequence of key frame states $\{S_{1},\cdots,S_{g}\}$ associated with the construction process, we seek to learn a program that models the inductive nature of the concept of tower. Online, the agent must generalize inductively to following an instruction as “construct a tower of blue blocks of size ten". Once the concept of tower is acquired, it is added to the concept library. This aids in learning the notion of a staircase as sequence of towers of increasing height.
We formalize the notion of inductive spatial concepts and formulate the learning objective. Let $C_{1},\cdots,C_{|\mathcal{L}|}$ represent the topological order of concepts present in the concept library $\mathcal{L}$ as per their structural complexity. An inductive spatial concept $C_{k}$ of size $n$ at position $p$ can be expressed as:
@@
$h(C_{k},n,p)=\underbrace{h^{\lambda}(C_{k},n-1,pos(.))}_{\text{Induction (I)}}\circ\underbrace{\prod\limits_{l=1}^{L(C_{k})}h\big{(}C_{k^{\prime}_{l}},\mathtt{size}(.),pos(.)\big{)}}_{\text{Composition (C)}}\circ\underbrace{\prod\limits_{l=1}^{L^{\prime}(C_{k})}\eta_{\theta}^{l}\big{(}pos(.)\big{)}}_{\text{Base (B)}},$ (1)
where, $\lambda\in\{0,1\}$, $k^{\prime}<k$, $0\leq L(C_{k}),L^{\prime}(C_{k})\leq o(|\mathcal{L}|)$ and $\eta_{\theta}^{l}$ is the set of primitive actions resulting in placement of object at a desired spatial position. Here, $pos(.)=pos(C_{k},l,n,p)$ and $\mathtt{size}(.)=\mathtt{size}(C_{k},l,n)$ are functions that predict the size and position of the structure to be constructed. For example, constructing a tower of size $n$ at position $p$ i.e., $h(Tower,n,p)=h^{\lambda=1}(Tower,n-1,p)\,\circ\,OnTop_{\theta}(p+(n-1)d)$. Here, $\lambda=1$, $L(C_{k})=0$, $L^{\prime}(C_{k})=1$, $\eta^{1}=OnTop$, $pos(Tower,l=1,n,p)=p+(n-1)\cdot d$ and $d$ is the dimension of the object. Thus, a tower can be described using induction and a single primitive base function with zero compositions. On the other hand, enclosures can be described as a composition of two rows and two columns without induction or base components. The functional space of inductive concepts ($h$) leads to a hypothesis space $\mathcal{H}$ of associated neuro-symbolic programs. Each goal-reaching demonstration corresponds to a particular instantiation of a given inductive concept, i.e. $H(C_{k},n,p)$, where the $p$ comes from the sequence of frames, and $n$, $C_{k}$ comes from $\Lambda$.
@@
Given demonstration, we aim to learn a generic representation for the given concept, which is general for all n, p i.e. $h(C_{k},\cdot,\cdot)$. We model this as learning a particular $H=h(C_{k},\cdot,\cdot)$ where $H\in\mathcal{H}$. Given (few) demonstrations of a human constructing a spatial structure, concept learning can be formulated as the Bayesian posterior [lake2015human, shah2018bayesian, lpp], $\mathtt{P}_{\mathcal{H}}(H|\Lambda,S_{1}..S_{g})\propto\mathtt{P}(S_{1}..S_{g}|\Lambda,H)\cdot\mathtt{P}(H|\Lambda)$. Here, the likelihood term associates a candidate program, and the prior term regularizes the program space. The maximum a-posteriori estimate, representing the learnt program, is obtained by optimizing the following objective:
$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\left(\mathtt{Loss}(\{S_{1}..S_{g}\},H(\Lambda))\right)-\log\left(\mathtt{P}(H|\Lambda)\right)\right]$ (2)
@@
4 Technical Approach: Learning Inductive Spatial Concepts
Figure 2: Approach Overview. We learn a neuro-symbolic program for inductive spatial concepts factored as (a) Sketch (b) Plan (c) Generalize. The example above shows the progressive realization of a program for the concept of a tower acquired by observing a single demonstration of building a tower of size three and its corresponding natural language instruction, “construct a tower of height three using green die."
We address the problem of estimating a succinct generalized program, Eq.  2, modeling an inductive spatial concept modeling structures whose construction is observed in a human demonstration. Direct symbolic search in the space of programs is intractable (particularly due to looping constructs needed for modeling induction) but can explicitly reason over previously acquired concepts. Alternatively, neural methods attempting to predict action sequence to attain the assembly are challenged by continual setting where concepts can increase over time building on previously learnt ones but are resilient to noise. Our approach blends both approaches and factors the concept learning task as:
  • Sketch: From the natural language instruction ($\Lambda$), we extract a task sketch ($H_{S}^{*}$) using an LLM that provides the signature (concept name and instantiated arguments) of the concept to be learned. When grounded in the initial scene of the demonstration, the task sketch provides a particular instance of the concept demonstrated in the given demonstration.
  • Plan: MCTS-based search using the already learnt concepts that outputs the sequence of grounded actions, best explaining the given demonstration.

### v3

# Sketch-Plan-Generalize : Learning and Planning with Neuro-Symbolic Programmatic Representations for Inductive Spatial Concepts

@@ Abstract
Effective human-robot collaboration requires the ability to learn personalized concepts from a limited number of demonstrations, while exhibiting inductive generalization, hierarchical composition, and adaptability to novel constraints. Existing approaches that use code generation capabilities of pre-trained large (vision) language models as well as purely neural models show poor generalization to a-priori unseen complex concepts. Neuro-symbolic methods (Grand et al., 2023) offer a promising alternative by searching in program space, but face challenges in large program spaces due to the inability to effectively guide the search using demonstrations. Our key insight is to factor inductive concept learning as: (i) Sketch: detecting and inferring a coarse signature of a new concept (ii) Plan: performing an MCTS search over grounded action sequences guided by human demonstrations (iii) Generalize: abstracting out grounded plans as inductive programs. Our pipeline facilitates generalization and modular re-use, enabling continual concept learning. Our approach combines the benefits of code generation ability of large language models (LLMs) along with grounded neural representations, resulting in neuro-symbolic programs that show stronger inductive generalization on the task of constructing complex structures vis-á-vis LLM-only and purely neural approaches. Further, we demonstrate reasoning and planning capabilities with learned concepts for embodied instruction following.
Machine Learning, ICML
[columns=3, title=Alphabetical Index, intoc]
@@ 1 Introduction
For robots to collaborate effectively with humans, they must quickly learn personalised concepts from just a few demonstrations. This ability to form inductive representations of novel grounded concepts and apply them for tasks beyond training is a hallmark of human intelligence (Tenenbaum et al., 2011; Chollet, 2019). In the context of robot learning, such representations are characterized by four fundamental properties: (a) it should be grounded directly in human demonstrations, reflecting the human’s intent rather than relying sorely on prior world knowledge, (b) it must generalize beyond observed examples; e.g., from a few towers of certain heights, infer how to build towers of arbitrary heights, (c) it should allow for the hierarchical interpretation of increasingly complex concepts as compositions of simpler ones; e.g., a staircase as a sequence of towers of increasing height, (d) it should be easily modifiable to incorporate additional constraints in new instructions, such as ensuring no green block is placed near a blue or yellow one in a staircase.
Existing approaches fall short of fully satisfying these criteria. Recent advances in large language models (LLMs) have enabled general-purpose planners. These planners use world knowledge to identify relevant concepts and generate plans from language task specifications (Singh et al., 2022; Liang et al., 2023; Ahn et al., 2022; Achiam et al., 2023; Huang et al., 2022a). However, they often struggle to learn new concepts from demonstrations, especially when the concepts are linguistically novel. This is because they rely heavily on prior world knowledge, which limits their faithfulness to demonstrations and their ability to generalize inductively. Purely neural approaches, such as (Liu et al., 2023), can learn directly from demonstrations but typically fail to generalize beyond training data because they (a) do not explicitly model symbolic induction and (b) lack modularity, restricting their ability to reuse learned concepts. In contrast, neuro-symbolic approaches (Grand et al., 2023; Ellis et al., 2021) offer a more promising alternative. They leverage LLMs or enumerative program search to discover generalizable programs directly in the program (string) space. While these models outperform purely LLM-based or neural methods in terms of generalization, they still face significant challenges: enumerative searches become computationally infeasible in complex program spaces (e.g., Python programs), and such search cannot be easily guided by human demonstrations.
@@
Figure 1: Problem Overview. Our goal is to enable an embodied agent to learn grounded and generalizeable representations for spatial abstractions possessing a notion of induction (e.g., constructing a tower, row or their combinations such as staircases, boundary etc.). (Left) A human demonstrates the construction of a row and tower of size three. (Right) The agent learns program representation that enables inductive generalization to novel structures (varied sizes and visual attributes) and expresses complex concepts as hierarchical composition of previously acquired ones. E.g., tower as blocks placed one on top of another and a pyramid as rows of decreasing size.
In response, we present a neuro-symbolic agentic framework that (i) identifies when learning a new concept is required, (ii) requests human demonstrations to learn an inductively generalizable representation of the unknown concept, and (iii) uses the learned concepts to generate control policies for novel instruction-following tasks with discrete constraints. Unlike prior methods, our concept learning approach, called SPG, factorizes the concept learning task into three stages: (a) Sketch: Given a language-annotated demonstration of a novel concept, an LLM is used to postulate a function signature. (b) Plan: Using MCTS, grounded action sequences are discovered by maximizing a reward that measures how closely the constructed structure matches the human demonstration. (c) Generalize: The code-generation capability of an LLM is leveraged to distill grounded plans from multiple demonstrations into an inductively generalizable program. This enables learning of inductively generalizable representations while restricting the search to the grounded actions, which can be effectively guided through dense rewards from human demonstration and is less expensive than searching directly in the program space. We also show how the learned concepts can be used for goal-conditioned planning in novel input scenes and instructions, even when visual or spatial constraints are present. Our experiments demonstrate that our framework can accurately learn both simple and complex concepts from just a few demonstrations across various spatial structures. Moreover, it exhibits strong inductive generalization in out-of-distribution settings, significantly outperforming baseline methods.
@@ 2 Related Works
Concept Learning: The problem of acquiring higher-order programmatic constructs is often modelled as Bayesian inference over a latent symbol space given observed instances. Seminal works have demonstrated efficient inference over latent generative programs to express handwritten digits (Lake et al., 2015), 2D drawings (Ellis et al., 2018), grid-world puzzles (Wang et al., 2024a), visual question answering (Mao et al., 2019), goal-directed policies (Silver et al., 2019), and compressed or refactored code (Grand et al., 2023; Ellis et al., 2021). These approaches typically either (i) search program space using a probabilistic grammar, which is infeasible for large program spaces like Python, or (ii) use LLMs to directly infer programs, which may produce physically ungrounded outputs. Furthermore, because partially generated programs cannot be evaluated, direct search in program space does not effectively leverage demonstrations to guide the search. In contrast, our approach first searches for grounded action sequences that can be efficiently guided by dense rewards from human demonstrations. We then use an LLM to abstract these sequences into a program, making program inference more tractable and efficient. While prior works focus on 2D problems without considering physical plausibility (for example, towers can be drawn from top to bottom but cannot be constructed that way in reality), we expose the search to assessing the physical construction plausibility, thereby learning physically grounded concepts.
Learning-to-plan Methods: Our work is complementary to efforts that learn symbolic constructs for efficient planning. Works such as (Silver et al., 2023, 2024; Liu et al., 2024), infer state-action abstractions for planning by querying large pre-trained models or by optimizing a goal attainability objective. This paper, instead, focuses on learning a representation for complex spatial assemblies as inductive programs leading to the ability to infer complex goal specifications which can then be combined with aforementioned works for synthesize efficient plans to realize complex assemblies. Works such as (Li et al., 2020) learn to construct structures by encoding relational knowledge via graph neural network. However, this effort suffers from poor generalization to unseen examples, (e.g., tower of larger size) and do not possess a mechanism to re-use previously acquired concepts. Works such as (Wang et al., 2023a, 2024b) shows lifelong learning of skills by learning to plan high-level tasks through composition of simple skills for simulated agents. Others (Wan et al., 2023; Parakh et al., 2023) initiate new skill acquisition upon detecting task failure, building a library of skills over time. However, they do not model inductive use of learned concepts and initiate skill acquisition only upon failure as opposed to learning continually even from goal-reaching demonstrations.
Robot Instruction Following: Instruction following involves grounding symbolic constructs expressed in language with aspects of the state-action space such as object assemblies (Paul et al., 2018; Collins et al., 2024; Lachmy et al., 2022), spatial relations  (Tellex et al., 2011; Kim et al., 2024), reward functions  (Boularias et al., 2015), or motion constraints  (Howard et al., 2014). These works assume the presence of grounded representation for symbolic concepts and only learn associations between language and concepts. In contrast our work jointly learns higher-order concepts composed of simpler concepts along with their grounding in the robot’s state and action space. Others (Singh et al., 2022; Wang et al., 2023b; Ahn et al., 2022; Liang et al., 2023) leverage prior-knowledge embodied in large vision-language models to directly translate high-level tasks to robot control programs. Our experiments (reported subsequently) demonstrate their limitation in outputting programs for structure assembly-type tasks that require long-range (inductive) spatial reasoning and consideration of physical plausability of construction. Our approach addresses this problem by coupling abstract task knowledge from pre-trained models with physical reasoning in the space of executable plans.
3 Preliminaries and Problem Setting
@@
We consider an embodied agent that uses a visual and depth sensor to observe its environment and can grasp and release objects at specified poses. We represent the robot’s domain as a goal-conditioned MDP $<\mathcal{S},\mathcal{A},\mathcal{T},g,\mathcal{R},\gamma>$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ is the transition function, $g$ is the goal, $R$ is the reward model and $\gamma$ is the discount factor. The agent’s objective is to learn a policy that generates a sequence of actions from an initial state $s_{0}$ to achieve the goal $g$ in response to an instruction $\Lambda$ specifying the intended goal from a human. We assume that the agent possesses a model of semantic relations (e.g., left(), right() etc.) as well as semantic actions such as moving an object by grasping and releasing at a target location. Such modular and composable notions can be acquired from demonstrations via approaches outlined in (Kalithasan et al., 2023; Mao et al., 2019, 2022). Such notions populate a library of concepts $\mathcal{L}$ available as grounded executable function calls. Following recent efforts (Liang et al., 2023; Huang et al., 2022b; Ahn et al., 2022; Singh et al., 2022) in representing robot control directly as executable programs, we represent action sequence corresponding to a plan as a program consisting of function calls to executable actions and grounded spatial reasoning.
@@
Our goal is to enable a robot to interpret and learn the concepts in instructions such as “construct a tower with red blocks of height five". Specifically, we aim to learn spatial constructs like a tower that requires sequential actions that repeatedly place a block on top of a previously constructed assembly, a process akin to induction. Given a few demonstrations of constructing a spatial assembly, $\mathcal{D}$, each consisting of natural language description $\Lambda$ (“construct a tower of red blocks of size five") and a sequence of key frame states $\{S_{1},\cdots,S_{g}\}$ associated with the construction process, we seek to learn a program that models the inductive nature of the concept of tower. This learned representation should enable the agent to generalize inductively to new instructions, such as "construct a tower of blue blocks of height ten." Moreover, the learned concepts should facilitate the learning of more complex structures, which are challenging to represent using primitive actions alone. For example, the concept of a "tower" should assist in learning a "staircase," which can be represented as a sequence of towers of increasing heights.
4 Representing Inductive Spatial Concepts
We formalize the notion of inductive spatial concepts and formulate the learning objective.
Inductive Spatial Concepts: A spatial structure is an inductive concept if its construction can be described recursively using a similar structure of smaller size or as a composition of other simpler structures. Formally, let $C_{1},\cdots,C_{|\mathcal{L}|}$ represent the concepts in the concept library $\mathcal{L}$. We define a partial order on $\mathcal{L}$ where a concept $C$ is "dependent on" $\widetilde{C}$ if $\widetilde{C}$ is a substructure of $C$. For example, a staircase is dependent on a tower, and $X$ (cross) is dependent on diagonals, and so on. This partial order is referred to as structural complexity, where a concept $C$ is more structurally complex than $\widetilde{C}$ if $C$ is dependent on $\widetilde{C}$. Without loss of generality, assume that $C_{1},\cdots,C_{|\mathcal{L}|}$ are written in topological order as per their structural complexity. Now, the construction of an inductive spatial concept $C_{k}$ of size $n$ at position $p$, denoted by $h(C_{k},n,p)$, is defined recursively as:
@@ 4 Representing Inductive Spatial Concepts
$h(C_{k},n,p)=\underbrace{h^{\lambda}(C_{k},n-1,pos(.))}_{\text{Induction (I)}}\circ\\ \underbrace{\prod\limits_{l=1}^{L(C_{k})}h\big{(}C_{k^{\prime}_{l}},\mathtt{size}(.),pos(.)\big{)}}_{\text{Composition (C)}}\circ\underbrace{\prod\limits_{l=1}^{L^{\prime}(C_{k})}\eta_{\theta}^{l}\big{(}pos(.)\big{)}}_{\text{Base (B)}}$ (1)
where, $\lambda\in\{0,1\}$, $k^{\prime}<k$, $0\leq L(C_{k}),L^{\prime}(C_{k})\leq o(|\mathcal{L}|),pos(.)=pos(C_{k},l,n,p)$ and $\mathtt{size}(.)=\mathtt{size}(C_{k},l,n)$ are functions that predict the size and position of the structure to be constructed.
  1. 1.
    Induction term: The first expression (I) represents induction vis-á-vis the possibility of constructing $C_{k}$ of size $n$ using $C_{k}$ of size $n-1$. Here, $\lambda$ depicts the absence($\lambda=0$) or presence($\lambda=1$) of the (I) term.
  2. 2.
    Composition term: The second term (C) expresses the construction of $C_{k}$ as a composition of prev. known concepts in the library ($\mathcal{L}$). The number of required compositions ($L(C_{k})$) depends on $C_{k}$ and $\mathtt{len}(\mathcal{L})$.
  3. 3.
    Base term: The third term (B) defines the base case where the construction of concept $C_{k}$ may include $L^{\prime}$ number of primitive actions. Here, $\eta_{\theta}$ are pre-trained primitive actions resulting in the repositioning of an object to the desired spatial position. For ex., a tower of size $n$ can be constructed by a (primitive) move-top action after constructing a tower of size $n-1$.
@@
Learning Objective: The functional space of inductive concepts ($h$) leads to a hypothesis space $\mathcal{H}$ of associated neuro-symbolic programs. Each goal-reaching demonstration corresponds to a particular instantiation of a given inductive concept, i.e. $h(C_{k},n,p)$, where the $p$ comes from the sequence of frames, and $n$, $C_{k}$ comes from $\Lambda$. We aim to learn a generic representation $H=h(C_{k},\cdot,\cdot)\in\mathcal{H}$ for the given concept, which is general for all $n$ and $p$. Given (few) demonstrations of a human constructing a spatial structure, concept learning can be formulated as the Bayesian posterior (Lake et al., 2015; Shah et al., 2018; Silver et al., 2019), $\mathtt{P}_{\mathcal{H}}(H|\Lambda,S_{1}..S_{g})\propto\mathtt{P}(S_{1}..S_{g}|\Lambda,H)\cdot\mathtt{P}(H|\Lambda)$. Here, the likelihood term associates a candidate program, and the prior term regularizes the program space. The maximum a-posteriori estimate, representing the learnt program, is obtained by optimizing the following objective:
$H^{*}=\arg\underset{H\in\mathcal{H}}{\mathrm{min}}[\mathtt{Loss}(\{S_{1}..S_{g}\},Exec(H,\Lambda,S_{1}))\\ -\log\mathtt{P}(H|\Lambda)]$ (2)
@@
5 Learning Inductive Concepts from Demonstrations
Figure 2: Method Overview. We learn a neuro-symbolic program for inductive spatial concepts factored as (a) Sketch (b) Plan (c) Generalize. The example above shows the progressive realization of a program for the concept of a staircase acquired by observing a single human demonstration of building a staircase of size four, and its corresponding language instruction
For estimating a succinct generalized program as per Eq.  2, a direct symbolic search in the space of programs can explicitly reason over previously acquired concepts, it is intractable, particularly due to looping constructs needed for modeling induction. Alternatively, neural methods that attempt to directly predict action sequence to explain the demonstrations are resilient to noise, but are challenged in the continual setting as the number of concepts can increase prohibitively over time. Our approach blends both approaches and factors the concept learning task as:
  • Sketch: From the natural language instruction ($\Lambda$), we extract a task sketch ($H_{S}^{*}$) using an LLM that provides the signature (concept name and instantiated arguments) of the concept to be learned.
  • Plan: An MCTS-based search is performed using the library concepts (learnt previously) which outputs a sequence of grounded actions that best explain the given demonstration.